Janny opened a savings account and deposited$300.00 as principal.The account earns 6% interest, compounded annually.What is the balance after 6 years?

1
Expert's answer
2021-12-21T07:07:04-0500

"A=P(1+\frac{r}{n})^{nt}"

Compounded annually n=1


"=\$300(1+\frac{0.06}{1})^{1×6}"


"=300(1.06)^6 =\$425.5557"

The balance after 6 years is $425.5557